OpenStudy (anonymous):

Question: While sitting on a tree branch 8.0 m above the ground, you drop a chestnut. When the chestnut has fallen 2.6 m, you throw a second chestnut straight down. What initial speed must you give the second chestnut if they are both to reach the ground at the same time? The answer I got but it isnt correct: 1st chestnut: X=X0+0.5at^2+V0t 5.4m= 0.5(9.81m/s^2)t^2 t=1.049s Chestnut 2: Used same equation 8m=0.5(9.81m/s^2)(1.049^2)+V0(1.049) V0=2.48sec If anyone can point out what I am doing wrong in this problem I would really appreciate the help.

OpenStudy (irishboy123):

1st chestnut: X=X0+0.5at^2+V0t 5.4m= 0.5(9.81m/s^2)t^2 here, you have calc'd time for 1st to fall 5.4, ie 8 - 2.6., from zero initial velocity but by time it has fallen 2.6m it already has a velocity

OpenStudy (anonymous):

so i use the X=X0+0.5at^2+V0t formula using the X=2.6 to determine the V0 then use that V0 and put that into the 5.4m= 0.5(9.81m/s^2)t^2+V0t equation?

OpenStudy (anonymous):

or nvm i would use V^2 = V0^2 +2a(2.6)

OpenStudy (anonymous):

I got final answer of V0 for chestnut two to be 7.14 m/s. Is that correct?